# TwinMarket: A Scalable Behavioral and Social Simulation for Financial Markets

### v2

@@ 1 Introduction
Recently, large language models (LLMs) have demonstrated the ability to capture intricate patterns in language use, decision-making, and social interactions [1, 2]. These capabilities have positioned LLMs as powerful tools for agent-based simulations in human behavior studies across various social science fields, including conversational dynamics [3, 4, 5], emotional responses [6], and creative reasoning [7, 8]. Such advancements open new avenues for social science research, enabling the modeling of fine-grained individual behaviors at the micro-level and the simulation of emergent phenomena at the macro-level.
@@
The study of social emergence phenomena, where micro-level individual behaviors aggregate to form macro-level collective outcomes, has long been a central focus in social science [13, 14]. However, current methods often struggle to capture the dynamic interplay between individual biases (e.g., overconfidence) and emergent market phenomena (e.g., volatility clustering) [15, 16]. Recent advances in LLMs offer a promising avenue to address this challenge, enabling the simulation of adaptive agents whose decisions are shaped by both private information and social interactions. Motivated by this potential, we propose TwinMarket, a novel multi-agent framework that leverages LLMs to simulate investor behavior in a stock market environment. By modeling agents that independently make investment decisions while interacting through a simulated social media platform, TwinMarket provides a unified perspective on how micro-level behaviors drive macro-level market dynamics.
As illustrated in Figure 1, a key innovation of TwinMarket is its use of the Belief-Desire-Intention (BDI) framework [17] to structure and visualize the cognitive processes of agents, providing a transparent and interpretable modeling approach. Additionally, we develop a social environment that enables agents to engage in dynamic information exchange and social influence processes, enabling the study of phenomena such as the emergence of opinion leaders and the dynamics of information cascades. Finally, by scaling our simulations to multiple financial assets and larger populations, we investigate how group size and interaction complexity affect collective behavior and emergent phenomena.
Our research makes three key contributions: (1) Real-World Alignment: Our framework is grounded in established behavioral theories and calibrated with real-world data, ensuring theoretical rigor and realistic human behavior modeling. This enables the study of emergent social phenomena. (2) Dynamic Interaction Modeling: By leveraging an LLM-based framework, we capture diverse human behaviors and their interactions, particularly in the context of information propagation. This allows us to model how opinions spread, influence decision-making, and dynamically shape collective market behaviors in real time. (3) Scalable Market Simulations: We conduct large-scale simulations to analyze the impact of group size on market behavior, providing new insights into price fluctuations and emergent dynamics.
@@ 2 Background
Traditional Agent-Based Models face two major challenges in simulating complex social systems: (1) customizing sophisticated economic decision-making for individual agents and (2) enabling agents to actively perceive and respond to external environmental changes. Recent research has made progress by incorporating LLMs into agent-based simulations. CompeteAI [18] introduces a framework for modeling competitive dynamics; however, its agents rely on predefined prompts, limiting their ability to simulate nuanced economic behavior. EconAgent [19] designs economic agents with perception, reflection, and decision-making capabilities to model macroeconomic dynamics but overlooks the influence of multi-agent interactions on emergent macroeconomic phenomena. Similarly, ASFM [20], which models stock market responses to regulatory policies, fails to adequately account for the role of agent interactions and collective decision-making in shaping market trends. Table1 provides a comparison of these systems with TwinMarket, highlighting the differences in agent customization, social interactions, and simulation scale.
System Agent Customization Social Interactions Simulation Scale
CompeteAI Simple predefined prompts Competition between 2 restaurants Less than 100 agents
EconAgent Uses demographic info No interaction considered 100 agents
ASFM Two fixed strategies No interaction considered Not disclosed
TwinMarket Highly customizable agents Rich multi-agent interactions 1000 agents, scalable
Table 1: Comparison of TwinMarket with Existing Literature
@@
3 Framework of TwinMarket
At the micro level, we focus on realistically modeling individual user behavior in the stock market, capturing decision-making processes and interactions grounded in decision theories and validated by real-world trading patterns. Moving to the macro level, we construct an interactive, forum-like environment where distributed agents can exchange information, form opinions, and influence each other’s decisions, thereby simulating the emergent properties of real-world social systems.
3.1 Data Sources and Simulation Setup
Usage Source Quantity/Time Period Data Type
Construct Initial User Profile Xueqiu1 639 users User Profile
11,965 transactions Transactions Details
Train Stock Recommendation System Guba Stock Forum2 83,246 transactions Transactions Details
@@ 3.1 Data Sources and Simulation Setup
Build Information Retrieval Database Sina5  & 10jqka6 Jan 2023 - Dec 2023 News Articles
@@
Table 2: Summary of data sources for TwinMarket.
Data Sources.
As shown in Table 2, TwinMarket integrates real user profiles, transacation details, stock data, news and announcements to create a realistic social simulation environment. Details about our data sources can be found in Appendix B.
Initial User Profile Construction.
Using real transaction data and user profiles from Xueqiu, we computed behavioral biases following Sui’s pioneering work [26] and extracted key characteristics like demographics to construct a joint distribution. This distribution was then used to sample and generate distinct agents, each initialized with trading records to provide a historical trading baseline. Additionally, each agent was initialized with a unique belief about the market environment and future expectations.Figure 2 illustrates an example of a generated agent, with further details on the generation process provided in Appendix C.
Market Environment Setup.
TwinMarket simulates a dynamic trading environment where agents’ decisions drive stock prices through an order-driven matching system. Initialized with given trading records, agents collectively determine price movements once the simulation begins, ensuring a market environment consistent with real-world dynamics and grounded in financial principles. Technical details are discussed in Section 3.4.
3.2 Micro-Level Simulation: Individual Behaviors
To model individual behaviors in financial markets, we employ the Belief-Desire-Intention framework [17], which is widely recognized for its ability to replicate human decision-making processes. However, in many simulations, agents are often exposed to homogeneous information flows, leading to behavioral convergence and a lack of meaningful heterogeneity [27]. To address this limitation, our approach controls each agent’s perception field by tailoring the information accessible to them based on their unique profiles.
The BDI framework allows us to construct individual user agents that mimic the behaviors and chain-of-thought [8] of real investor in the stock market. Specifically, the BDI model involves three interconnected components: belief, which represents the agent’s evolving understanding of the market environment based on incoming data and predictions; desire, which captures the agent’s objectives or preferences, guiding its pursuit of specific information or outcomes; and intention, which defines the concrete actions or strategies that the agent executes to achieve its goals.
At the micro-level, our agents are equipped with three key functionalities: signal perception, planning, and decision-making. These capabilities allow them to replicate the nuanced behaviors of real-world investors. While financial decision-making may appear complex, it can be broken down into a sequence of smaller, observable actions. For example, individuals receiving market information such as stock price movements, breaking news, or social media discussions process this information to align it with their financial goals or aspirations (desire), form an understanding of the market (belief), and ultimately execute specific trading or interact with other users through social media (intention). By replicating these granular decision-making processes, the agents establish a solid foundation for simulating the broader social structures and dynamics of financial markets.
Each agent is assigned a unique role within the BDI framework, with clearly defined responsibilities and action spaces, as detailed in Table 3. A comprehensive description of how these agents operate within the simulation workflow will be presented in Section 3.4.
Table 3: The definition and action space for each BDI agent.
Agent Definition Action Space
Belief Market understanding & prediction -
Desire Information retrieval goals Query, Search
Intention Environment interaction & decision making Social Media, Trading
Figure 3: The overall workflow of the TwinMarket involves the following iterative process: ① User Data Initialization: User profiles are initialized using Snowball data and their calculated behavioral biases. ② Daily Simulation: Using LLMs to simulate daily user behavior, distinguishing between trading and non-trading days. ③ Iterative Trading & Update: The system iteratively executes trades on trading days and updates user profiles on all days, looping back to simulation. ④ Post-Simulation Validation: System performance is validated using micro and macro metrics after simulation completion.
3.3 Macro-Level Simulation: Social Interactions
In the financial market, information flow is a fundamental driver of collective behavior and market dynamics. Graphs provide a natural and powerful framework for modeling how information propagates through a network of interconnected agents, making them indispensable for macro-level simulations [28].
@@ 3.3 Macro-Level Simulation: Social Interactions
The market we are designing is inherently dynamic, with agents whose behaviors evolve over time. This dynamic nature requires the system to be adaptive and scalable, capable of responding to shifts in individual agent behavior and interactions. To meet these requirements, we adhere to two fundamental principles. First, the graph structure must be dynamic, able to evolve in response to the changing patterns of individual agents. Second, the agents must be capable of aggregating and processing environmental information through their interconnected relationships. These principles guide the construction of a framework that links micro-level interactions with macro-level systemic dynamics, enabling the system to adapt and scale effectively.
Social Network Construction.
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be a social network where $\mathcal{V}=\{u_{1},u_{2},\dots,u_{n}\}$ represents the set of user nodes, and $\mathcal{E}$ represents the set of edges connecting users based on their trading behavior. Each user trades stock indexes belonging to industries $I=\{i_{1},i_{2},\dots,i_{m}\}$, as defined in Appendix B.1. We define the graph such that users with more similar trading behaviors have stronger connections in the social network. This reflects the idea that individuals with similar trading patterns are more likely to share information and influence each other, leading to higher homogeneity in the information they receive.
To model the dynamic nature of trading behavior (detailed in Appendix D.1), we introduce a time decay factor $\lambda$. This factor ensures that more recent trades have a greater influence on the graph structure, while older trades contribute less. For a trading record $T(u,i,t)$, where user $u\in\mathcal{V}$ trades in industry $i\in I$ at time $t$, we compute the time-decayed weight of their interactions as:
$w(u,i)=\sum_{t\in T(u,i)}e^{-\lambda\cdot\Delta t}$
where $\Delta t$ is the time difference (in days) between the current time and $t$. To quantify similarity, we calculate the weighted Jaccard similarity $S(u_{1},u_{2})$ between users $u_{1}$ and $u_{2}$:
$S(u_{1},u_{2})=\frac{\sum_{i\in I}\min(w(u_{1},i),w(u_{2},i))}{\sum_{i\in I}\max(w(u_{1},i),w(u_{2},i))}$
An edge $(u_{1},u_{2})\in\mathcal{E}$ is created with a weight $S(u_{1},u_{2})$, representing the strength of the connection between the two users.
Information Aggregation Mechanism.
To control the perception field of each agent and ensure they receive relevant information, we design an information aggregation mechanism based on social network interactions. For a target user $u_{t}\in\mathcal{V}$, let $\mathcal{N}(u_{t})\subseteq\mathcal{V}$ denote the set of neighboring users of $u_{t}$ in the social network $\mathcal{G}$, and let $S(u_{t},u)$ represent the similarity between $u_{t}$ and a neighboring user $u\in\mathcal{N}(u_{t})$. We consider only those neighboring users whose similarity with $u_{t}$ exceeds a threshold $\tau$, i.e., $S(u_{t},u)>\tau$.
For each such user $u$, we retrieve their posts $P(u)$ created within a specified time range $[t_{s},t_{e}]$, where $t_{s}$ and $t_{e}$ are the start and end timestamps, respectively. Each post $p\in P(u)$ is associated with metadata, including its creation time $t_{p}$, upvotes $u_{p}$, and downvotes $d_{p}$.
To rank the posts, we compute a hot score $h(p)$ for each post $p$ using a logarithmic scaling of net votes combined with a time-decay factor:
$h(p)=\begin{cases}\frac{\log_{10}(u_{p}-d_{p}+1)}{(T_{p}+1)^{1.8}},&\text{if }u_{p}-d_{p}>0\\ -T_{p},&\text{otherwise}\end{cases}$
where $T_{p}=\max\left(\frac{t_{c}-t_{p}}{86400},0.1\right)$ represents the time elapsed in days since the post was created, with $t_{c}$ being the current timestamp. This formulation ensures that highly engaged posts remain visible while older or negatively received posts decay more rapidly.
The posts are then sorted in descending order of their hot scores, and the top $k$ posts are recommended to the target user $u_{t}$, where $k$ is a predefined maximum number of recommendations. This mechanism leverages the social network structure to aggregate information from users with significant similarity to $u_{t}$, while prioritizing content based on engagement and recency.
3.4 Workflow
Our design philosophy, as illustrated in Figure 3, simulates the complete workflow of users in real-world trading environment, from acquiring external information and making trading decisions to sharing insights with others.
User Behavior.
Users start by engaging with curated posts, where they can upvote, downvote, or repost content that resonates with their perspectives. Following this, they receive timely environmental updates, such as news and market developments, to stay well-informed. These inputs, combined with their current beliefs, shape users’ desires to seek specific information through targeted queries across news, announcements, and stock data. As users process this new information, they continuously update their beliefs, which in turn influence their evolving desires and ultimately form their intentions. These intentions lead to transaction decisions, while users actively share their beliefs and market views through posts, enriching the system’s collective knowledge base for future decision-making processes.
Dynamic Environment.
To enable personalized decision-making and ensure distinct behavioral patterns across personas, we regulate each user’s perception field using recommendation algorithms. The stock recommendation system is driven by users’ portfolios and real trading data. Meanwhile, the post recommendation system dynamically adjusts based on post scores, which fluctuate in real-time and influence their ranking and visibility. Additionally, each user interacts with these recommendations selectively, guided by their individual perception field. Once a user interacts, information propagates dynamically, forming post-propagation chains that circulate throughout the market. Before trading, users can search for relevant news and announcements (Table 2) and access technical indicators (e.g., trading volumes) and fundamental indicators (e.g., P/E ratios). To maintain valuation consistency, fundamental metrics are scaled using a ratio that aligns simulated prices with real-world values [^1] Since fundamental metrics cannot be solely derived from agent-driven trading, they are adjusted based on the relative difference between simulated and real-world prices..
4 Results and Analysis
We conduct experiments on a network of 100 users, starting with a micro-level analysis to examine individual decision-making differences across various scenarios. We then shift to a macro-level perspective, evaluating how collective behaviors influence overall market dynamics. Our findings highlight a key aspect: the dynamic interplay between information propagation and collective behavior, where self-fulfilling prophecies, rumors, and shifts in sentiment trigger market fluctuations. Specifically, we observe how these phenomena drive trading behaviors and impact market stability, phenomena that were previously difficult to capture using traditional ABMs.
4.1 Experimental Settings
Based on our framework, we implemented the environment as a stock market with 10 representative indexes22210 indexes are aggregated from SSE 50, detailed in Appendix B.1. and 100 market participants entirely powered by GPT-4o [29]. The participants, consisting of fundamental and technical analysts, operate in a simulation period of 5 months, from June 15, 2023, to November 15, 2023. Each agent is initialized with a unique set of beliefs based on a predefined persona (detailed in Appendix C.5), which influences their behavior bias and market assessment. The agents develop daily trading strategies and can freely choose to buy, sell, or hold positions, rather than being forced to trade. To facilitate information exchange, we implemented a stock forum where agents can share insights and market sentiment. The simulation will terminate early if any stock price exceeds predefined volatility limits or if there is a significant decrease in market participation.
4.2 Micro-analysis: Self-fulfilling Prophecy

### v3

# Kaidi Zhang1, Yunmiao Zhang2, Honghai Yu2, Yan Hu1, Benyou Wang1 1

@@ 1 Introduction
Recent advances in large language models (LLMs) have showcased their remarkable capacity to capture complex patterns in language, decision-making, and social interactions [1, 2]. These capabilities have positioned LLMs as powerful instruments for agent-based simulations in the study of human behavior, supporting a wide range of social science applications—including conversational dynamics [3, 4, 5], emotional responses [6], and creative reasoning [7, 8]. These developments open new frontiers for social science research, enabling the modeling of fine-grained individual behavior at the micro-level and the simulation of complex, emergent phenomena at the macro-level.
@@
The study of social emergence phenomena, where micro-level individual behaviors aggregate to form macro-level collective outcomes, has long been a central focus in social science [13, 14]. However, current methods often struggle to capture the dynamic interplay between individual biases (e.g., overconfidence) and emergent market phenomena (e.g., volatility clustering) [15, 16]. Recent advances in LLMs offer a promising avenue to address this challenge, enabling the simulation of adaptive agents whose decisions are shaped by both private information and social interactions. Motivated by this potential, we propose TwinMarket, a novel multi-agent framework that leverages LLMs to simulate investor behavior in a stock market environment. By modeling investors that independently make investment decisions while interacting through a simulated social media platform, TwinMarket provides a unified perspective on how micro-level behaviors drive macro-level market dynamics.
The key innovation of TwinMarket is its use of the Belief-Desire-Intention (BDI) framework [17] to structure and visualize the cognitive processes of agents, providing a transparent and interpretable modeling approach. Additionally, we develop a social environment that enables agents to engage in dynamic information exchange and social influence processes, enabling the study of phenomena such as the emergence of opinion leaders and the dynamics of information cascades. Finally, by scaling our simulations to multiple financial assets and larger populations, we investigate how group size and interaction complexity affect collective behavior and emergent phenomena.
Our research makes three key contributions: (1) Real-World Alignment: Our framework is grounded in established behavioral theories and calibrated with real-world data, ensuring theoretical rigor and realistic human behavior modeling. This enables us to accurately reproduce certain social phenomena, allowing us to further study the mechanisms behind these phenomena. (2) Dynamic Interaction Modeling: By leveraging an LLM-based framework, we capture diverse human behaviors and their interactions, particularly in the context of information propagation. This allows us to model how opinions spread, influence decision-making, and dynamically shape collective market behaviors in real time. (3) Scalable Market Simulations: We conduct large-scale simulations to analyze the impact of group size on market behavior, providing new insights into price fluctuations and emergent dynamics.
@@ 2 Background
Traditional ABMs face two major challenges in simulating complex social systems [18]: (1) customizing sophisticated economic decision-making for individual agents and (2) enabling agents to actively perceive and respond to external environmental changes. Recent research has made progress by incorporating LLMs into agent-based simulations. CompeteAI [19] introduces a framework for modeling competitive dynamics; however, its agents rely on predefined prompts, limiting their ability to simulate nuanced economic behavior. EconAgent [20] designs economic agents with perception, reflection, and decision-making capabilities to model macroeconomic dynamics but overlooks the influence of multi-agent interactions on emergent macroeconomic phenomena. Similarly, ASFM [21], which models stock market responses to regulatory policies, fails to adequately account for the role of agent interactions and collective decision-making in shaping market trends.
Table 1: Existing financial agent systems.
System Agent Design Scale / Interaction
CompeteAI Fixed prompts 2 agents, competitive
EconAgent Demographic-based 100 agents, no interaction
ASFM Two strategies N/A, no interaction
TwinMarket Fully configurable 1000+, rich interaction
@@
3 TwinMarket
3.1 Overall Design of TwinMarket
The design philosophy of TwinMarket emphasizes both micro-level and macro-level simulation. At the micro-level, we focus on modeling user behavior, as illustrated in Figure 2(a). This entire process is driven by BDI framework, which we will discuss in Section 3.2. At the macro-level, as illustrated in Figure 2(b), we simulate two core infrastructures of real-world financial ecosystem: an order-driven trading system that settles transactions and updates the stock index accordingly (detailed in Appendix G.4), and a social media environment that personalizes information delivery based on user behavior while enabling social interaction among agents.
Figure 2: The simulation workflow driven by the Belief-Desire-Intention framework. The micro-level simulation focuses on modeling individual user behavior, while the macro-level simulation addresses the dynamics of the social media platform and trading system.
Additionally, TwinMarket is a data-driven platform that leverages real-world data sources to construct a realistic and credible simulation environment, as summarized in Table 2. Our synthetic user data is sampled based on behavioral patterns observed from real users, and enriched with actual stock data, news, and company announcements. This setup enables simulated users to make informed decisions, allowing us to observe the emergence of macro-level market phenomena from individual interactions.
Table 2: Summary of data sources for TwinMarket. Further details about our data sources are provided in Appendix E.
Usage Source Quantity or Time Period Data Type
User Profile Initialization Xueqiu1 639 users User Profile
11,965 transactions Transaction Details
Stock Recommendation Guba Stock Forum2 83,246 transactions Transaction Details
@@ 3.1 Overall Design of TwinMarket
World Knowledge Database Sina5  & 10jqka6 Jan 2023 - Dec 2023 News Articles
@@
3.2 Micro-Level Simulation: BDI-driven User Modeling
In user modeling, the BDI framework plays a central role in simulating the cognitive processes of users. It enables each agent to perceive the market environment, integrate that information with its individual profile, and form a unique chain of reasoning that guides decision-making. Figure 2(a) shows how an LLM-based agent operates in a daily loop: receiving news and posts, querying from knowledge base, making trading decisions, and interacting via social platforms. Each action in this loop is governed by the agent’s beliefs, desires, and intentions, enabling nuanced and individualized behavioral responses.
Belief. Belief provides the dynamic foundation for how an agent perceives and interprets its environment. To ensure agents follow consistent personas, we inject system prompts that encode individualized traits. Using real transaction data and user profiles from Table 2, we first estimate behavioral biases and then extract key characteristics such as demographics and investment styles. These features are used to synthesize user profiles, forming the basis of our agent society. The distribution of these synthetic profiles closely aligns with findings [^1] Order-level transaction data is sensitive, and our goal is to show that the data we collect is reliable, especially compared to prior studies with access to proprietary firm data. from prior financial research [27]. As user profiles are initialized, beliefs encoded as system prompts guide agents to read and interpret news and social media posts. These perceived signals, grounded in agent traits, shape their evolving understanding of the environment and serve as the cognitive basis for subsequent decision-making.
After completing all actions for the day, each agent evaluates its behavior based on both its actions and the environmental feedback it received. Through this process, the agent dynamically updates its beliefs, reflecting an evolving understanding and interpretation of the market. This captures the notion that a user’s perception and judgment of the environment are not static but shaped continuously by experience and new information. To track belief dynamics and explore micro-to-macro linkages, agents output a sentiment score [^2] The sentiment score ranges from 1 to 5 and is computed as the average of five self-assessed dimensions, based on user’s views on: economic fundamentals, market valuation levels, short-term market trends, the sentiment of surrounding investors, and self-confidence, as detailed in Appendix G.1. at each update step.
Desires. Desires represent the goals or objectives an agent aims to achieve based on its beliefs. While beliefs provide an understanding of the environment, agents autonomously generate queries to retrieve relevant stock and news information, enhancing their comprehension of the market context.
Intentions. Intentions represent the committed plans or actions that an agent chooses to pursue among its possible desires. While desires reflect what an agent wants, intentions define what it actively decides to do, given its current beliefs and constraints. These intentions drive the agent to make final trading decisions and to express or amplify opinions through posting or reposting content on social media.
3.3 Macro-Level Simulation: Dynamic Social Network
@@
The relationship between micro-level agent behavior and macro-level market dynamics is central to understanding complex financial systems [28]. The market we design is inherently dynamic, driven by agents whose behaviors evolve over time. This requires the system to be both adaptive and scalable, capable of responding to shifts in individual behavior and interaction patterns. To achieve this, we follow two key principles. First, the graph structure must evolve dynamically to reflect changes in agent connectivity. Second, agents must be able to aggregate and interpret environmental information through their networked relationships.
Constructing Dynamic Social Network. We model the social network as a dynamic graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ is the set of users and $\mathcal{E}$ denotes weighted edges that capture similarity in trading behavior. Each user trades industry-level stock index $I=\{i_{1},\dots,i_{m}\}$, and we assume that users with more similar trading patterns are more likely to interact. To reflect temporal dynamics, we apply exponential time decay to each trading record $T(u,i,t)$, resulting in a time-weighted trading intensity $w(u,i)=\sum_{t\in T(u,i)}e^{-\lambda\cdot\Delta t}$, where $\Delta t$ is the time elapsed since trade time $t$. We then compute the weighted Jaccard similarity [^3] We use a generalized Jaccard similarity for real-valued vectors, calculated by: $S=\sum_{i}\min(w_{1},w_{2})/\sum_{i}\max(w_{1},w_{2})$. between users to determine the strength of social connections  [29]. We provide a sensitivity analysis of graph construction parameters in Appendix D.1. Figure 3 shows the dynamic social network as a chord diagram. Colors indicate user groups with similar industry preferences (e.g., Manufacturing, Finance Service), and chord thickness reflects the strength of trading similarity between groups.
Figure 3: Chord diagram of user dynamic social network.
Information Aggregation Mechanism. To control each agent’s perception field and ensure exposure to relevant information, we design an aggregation mechanism based on the dynamic social network. For a target user $u_{t}\in\mathcal{V}$, we consider neighbors $u\in\mathcal{N}(u_{t})$ with similarity $S(u_{t},u)>\tau$. For each such user, we retrieve their posts $P(u)$ within a time window $[t_{s},t_{e}]$, where each post $p\in P(u)$ contains metadata such as timestamp $t_{p}$, upvotes $u_{p}$, and downvotes $d_{p}$. We compute a hot score $h(p)$ for each post to balance popularity and recency:
Figure 4: Hot score decay: Post influence typically lasts 2 weeks.
$h(p)=\frac{\log_{10}(u_{p}-d_{p}+1)}{(T_{p}+1)^{1.8}},\quad T_{p}=\max\left(\frac{t_{c}-t_{p}}{\alpha},\epsilon\right),$
where $t_{c}$ is the current time, $\epsilon$ prevents division by zero, and $\alpha$ is a time normalization constant controlling decay rate. As illustrated in Figure 4, we set $\alpha$ such that a post’s influence typically lasts around two weeks before its score significantly diminishes. Posts from all qualified neighbors are ranked by their hot scores, and each user is recommended the top $k$ posts per day. This mechanism enables personalized and dynamic information exposure shaped by social similarity and engagement dynamics. More technical details about information on social media are further discussed in Appendix G.2.
4 Experimental Validation
To evaluate the realism and reliability of our simulation framework, we conduct a series of experiments comparing the outcomes generated by TwinMarket with real-world financial and behavioral data. Our validations are performed both in micro and macro level, which are conducted on a simulated social network with 100 user agents, where each agent is powered by GPT-4o. The simulation runs from 2023-6-15, and lasts for five months.
4.1 Micro-level Validation
We validate the micro-level realism of TwinMarket by examining emergent behavioral and economic patterns, with qualitative examples discussed in Appendix D.5.
Figure 5: Rising Gini coefficient, indicating widening wealth inequality.
Inequality in the Stock Market. TwinMarket exhibits emergent patterns of wealth inequality, demonstrating its micro-level fidelity. As shown in Figure 5, the Gini coefficient displays a clear upward trend, accompanied by a growing wealth share for the top 10% of agents and a declining share for the bottom 50%. This self-organizing wealth divergence arises from the heterogeneity and interactions among agents, echoing established findings in the economics literature [30]. It highlights the model’s capacity to generate realistic socio-economic dynamics from its micro-level specifications.
Table 3: Agent performance vs. Trading frequency in TwinMarket
Performance Avg. Turnover Avg. Return
Top 10% 4.02% 6.65%
Bottom 50% 7.03% -10.52%
Trading Activity and Returns. Table 3 shows the top 10% of agents (ranked by return) achieved an average return of 6.65% with a modest turnover rate of 4.02%. In contrast, the bottom 50% showed a higher average turnover rate of 7.03% but suffered a substantial negative return of -10.52%. This pattern reflects a well-documented phenomenon among retail investors, where higher trading frequency even before accounting for transaction costs correlates with poorer performance [31]. Such behavioral consistency further reinforces the micro-level validity of TwinMarket.
4.2 Macro-level Validation
To assess TwinMarket’s ability to replicate realistic market dynamics, we compare its simulated outputs against historical data by examining four widely recognized stylized facts. These are empirical statistical regularities consistently observed in real-world financial markets, and are commonly used as benchmarks for validating the realism of market simulation models [32].
Stylized Fact I.
Fat-tailed distributions [33]: Big market moves happen more often than what a normal distribution would suggest. This means extreme ups and downs are not as rare as we might expect.
Figure 6 shows TwinMarket’s price change distribution (Log Return on its x-axis) mirroring real-world markets by exhibiting fat tails. While most daily price changes are small (forming its sharp central peak), large, unexpected surges or plunges (the heavy tails) occur more frequently than a normal distribution would predict, capturing actual market risk. This realism stems from our autonomous agents, which, based on BDI cognitive frameworks, process information heterogeneously, leading to diverse and potentially conflicting sentiments. These divergent sentiments then propagate through the integrated social network, becoming amplified and polarized, which in turn directly leads to extreme market outcomes.
Figure 6: Fat-tailed price changes: more extreme movements.
Stylized Fact II.
Leverage Effect [34]: When prices go down, the market often becomes more volatile afterward. This means bad news makes the market more nervous than good news.
Figure 7 shows TwinMarket capturing the leverage effect, where volatility responds asymmetrically. Specifically, initial price drops (x-axis Negative Returns) provoke larger subsequent drops (y-axis Lagged Negative Returns), thereby amplifying market volatility. This realistic asymmetry is a direct consequence of our LLM agents’ design; their BDI cognitive architectures are intentionally embedded with behavioral biases, such as loss aversion. This design makes them, akin to humans, inherently more sensitive to adverse signals. The integrated social media platform further amplifies the spread of negative sentiment, resulting in more widespread, synchronized, and potentially destabilizing collective trading behavior.
Figure 7: Initial price drops can trigger subsequent price drops.
Stylized Fact III.
Volume-Return Relationship [35]: Larger price changes usually happen when more people are trading, showing a positive link between trading volume and return size.
Figure 8 shows TwinMarket’s ability to replicate the volume-return relationship. Specifically, it shows that substantial increases in trading activity (x-axis Volume Change) are typically accompanied by significant market price fluctuations (y-axis Log Return). This co-movement is a key indicator, as in real markets, of collective behaviors such as herding among participants. In TwinMarket, the emergence of this co-movement provides compelling evidence demonstrates that our LLM agents, through their interactions, successfully generate collective behavioral patterns consistent with those observed in actual financial market dynamics.
Figure 8: Corr. between trading volume & price changes.
Stylized Fact IV.
Volatility Clustering [36]: When the market is volatile, it often stays volatile for a while. When it’s calm, it usually stays calm too. This means that big or small price changes tend to come together, not appear randomly.
Figure 9 shows TwinMarket successfully replicating volatility clustering: periods of large price swings (high Volatility on y-axis) tend to cluster together, as do periods of small swings (low Volatility). Both simulated and real volatility series clearly show these alternating groupings of high and low market turbulence over time. This emergent persistence arises as impactful market-moving events, processed via the BDI cognitive architecture, induce sustained heightened belief sensitivity. The social network then perpetuates these collective cognitive states by circulating related discussions or sentiments, prolonging periods of either erratic or stable trading.
Figure 9: Big price swings group together; small ones too.
ABM Baselines Comparison.
Table 4: Comparing Stylized Facts (SF) across Real Data, TwinMarket, and ABM baselines.
System SF I SF II SF III SF IV
Real Data 7.26 0.14 $p{<}0.01$ 0.95
TwinMarket 5.24 0.11 $p{<}0.01$ 0.89
ABM-HPM 4.47 0.05 $p{<}0.01$ 0.82
ABM-BH 4.99 0.19 $p{<}0.01$ 0.72
We compare TwinMarket against two transitional agent-based models: the HPM Model [37], which captures investor strategy adaptation through herding behavior, predispositions, and reactions to price misalignments; and the BH Model [38, 18], which incorporates hyperbolic discounting and stochastic volatility to reflect heterogeneous investor behavior. As summarized in Table 4, TwinMarket consistently generates outputs that align more closely with real-world market statistics across all four stylized facts. This demonstrates its superior ability to replicate fundamental financial dynamics, thereby validating the empirical robustness of our simulation framework. These results were further validated for reproducibility and stability, as detailed in Appendix D.3.
The four stylized facts (SF) in Table 4 are quantified as follows: (1) Kurtosis (SF I): Measures tail heaviness; values greater than 3 indicate fat-tailed return distributions. (2) Negative Return Autocorrelation (SF II): Captures asymmetric responses to negative shocks. (3) Volume Return Correlation (SF III): A significant $p$-value ($<0.01$) suggests stronger volume–volatility linkage. (4) GARCH Volatility Clustering (SF IV): The sum $\alpha+\beta\approx 1$ reflects persistent volatility, a hallmark of financial time series [39].
5 Emergence Analysis
Building upon the empirical validation of our simulation results, we have demonstrated that TwinMarket can generate behavior and market patterns consistent with real-world financial and social dynamics. To further investigate how micro-level agent behaviors give rise to macro-level phenomena, we conduct an extended set of experiments aimed at uncovering the mechanisms of social and market emergence. These analyses focus on the interplay between individual belief updates, local interactions, and global outcomes, offering insight into how localized decision rules and information asymmetries can lead to systemic patterns such as bubbles, polarization, and volatility.
5.1 Micro-level Emergence: Self-fulfilling Prophecy